# Partial identification and unmeasured confounding with multiple treatments and multiple outcomes

### v2

@@ 4.3.2 Identification conditions with negative controls
While the focus of our paper is generally on partial identification, we show here that under certain scenarios, negative controls can point identify the causal effect of interest. This further shows the potential benefit of negative controls and provides connections between our work and identification scenarios in the literature on double negative controls (Miao et al., 2018; Shi, Miao, Nelson & Tchetgen Tchetgen, 2020; Shi, Miao & Tchetgen Tchetgen, 2020; Hu et al., 2023).
@@
A proof of this result can be found in Appendix A.5. While these conditions may seem abstract, they do hold in some plausible scenarios. For example, scenario 1 holds automatically when we have $m$ negative control exposures, i.e. negative control pairs with the same outcome as the estimand of interest, but different exposure contrasts. Identification can also be obtained if additional assumptions are made, such as the existence of a so-called confounding bridge function (Miao et al., 2018), though we point readers to the literature on double negative controls for a discussion of such assumptions. Lastly, it is important to emphasize that within our framework, negative controls are still useful even if these conditions do not hold, as they can reduce the widths of the partial identification region substantially compared with the widths without negative controls.
@@ 5 Simulation study
Here, we present a brief simulation study to highlight the benefits of characterizing the joint partial identification region for multiple estimands. For simplicity, we focus on a single data set with sample size $n=10^{6}$ to effectively eliminate sampling variability and focus on the uncertainty caused by the potential presence of unmeasured confounders. For brevity, we leave explicit details of the data generation process to Appendix D. To summarize, we consider a problem with $k=10$ exposures and $q=6$ outcomes, with $m=3$ unmeasured confounders. The true effect sizes for all estimands are between -1 and 1, and there are a number of treatments with no effect on specific outcomes, leading to multiple potential negative controls. We use linear models throughout, which contains the true data generating mechanism. Our inferential goal is to estimate the effect of a particular exposure on a single outcome. We scaled the confounding and effect sizes so that the estimate obtained assuming no unmeasured confounding is 0, while the true value of the estimand is 1. To illustrate the different assumptions or constraints that one can place, we derive the partial identification region for this estimand under a variety of different assumptions, summarized below.
@@
Understanding the health effects of air pollution is a critically important problem in public health research. There is a vast literature showing detrimental impacts of air pollution on a variety of health outcomes (Cohen et al., 2017; Burnett et al., 2018; Manisalidis et al., 2020). Nearly all of the research establishing causal effects of air pollution on public health is observational, however, and relies on certain untestable assumptions. In some settings, a natural experiment is available that helps to alleviate issues stemming from unmeasured confounding (Rich, 2017), but in many cases it is assumed that the observed covariates contain all confounders of the exposure-outcome relationship. Due to the importance of this public health problem, and the impact these studies have on environmental regulatory policy, it is important to assess whether these findings are robust to unmeasured confounding bias. Here, we used the proposed methodology to estimate the causal effect of ozone and specific components of PM2.5 on a variety of health outcomes, while accounting for the possibility that these are subject to unmeasured confounding bias.
We use claims data from Medicare Fee-for-Service enrollees over the age of 65 that were enrolled in the years 2000 to 2016 obtained from the Centers of Medicare and Medicaid Services. For each enrollee we have information on their age, sex, race, and whether they are Medicaid eligible, which functions as a proxy for low socioeconomic status, as well as their residential zip code. In addition to these individual level characteristics, we have a number of area level characteristics that are available through the United States Census Bureau and the Center for Disease Control’s Behavioral Risk Factor Surveillance System. These include covariates such as average body mass index, education level, population density, smoking rates, median household income and house value, and the percentage of owner occupied housing. We also adjust for meteorological variables, such as temperature and precipitation rates from the National Climatic Data Center.
@@
In our analysis, we obtain $k=6$ different air pollution exposures from two distinct publicly available sources: we obtain estimates of ammonium (NH4), nitrates (NO3), and sulfate (SO4) levels on a ($0.01^{\circ}\times 0.01^{\circ}$) grid from the Atmospheric Composition Analysis Group (Van Donkelaar et al., 2019) and we obtain estimates of elemental carbon (EC), organic carbon (OC), and ozone (O3) on a 1km by 1km daily grid from the Socioeconomic Data and Applications Center (Di et al., 2021; Requia et al., 2021). These exposures are aggregated to the zip code level so that they can be linked to the residential zip code obtained from Medicare, and are additionally aggregated to the yearly level as we are examining long term exposure to pollution. We have $q=6$ outcomes that consist of zip code level hospitalization rates for anemia, COPD, lung cancer, stroke, and asthma which are obtained based on ICD9/10 diagnosis codes. Additionally, we have access to a negative control outcome, which is COPD rates in the year prior to exposure. Our unit of analysis is a United States zip code, and all covariates are aggregated up to the zip code level by taking their averages or proportions within each zip code. We standardize all exposures and outcomes to have mean zero and standard deviation one prior to running the analysis so that the effect estimates are all on the same scale across outcomes and exposure combinations. For example, an estimate of 0.1 for a particular shift in exposures would be interpreted as being expected to change the outcome by 0.1 outcome standard deviations. The estimands we look at are the effects of shifting one exposure from their 1st to 3rd quartile, while holding the remaining exposures fixed at their mean value. We estimate the effects of the exposures (and confounders) on the outcome using a linear regression model, and use the aforementioned approaches to assess the robustness of these estimates to unmeasured confounding bias. For more details on the estimation of all model parameters, including factor model parameters, see Appendix C.
@@
We also note that the shared confounding assumption (Assumption 4) is plausible in this application. All outcomes correspond to hospitalization rates for different diseases, and it is reasonable that certain characteristics, which are not measured in our observed covariates, are broadly associated with poor health outcomes. For instance, these unmeasured characteristics likely reflect features like health-seeking behavior or proximity to and/or quality of health care access. Similarly, exposures tend to be elevated in areas with various pollutant emitting sources such as roadways, power plants, and factories. Certain communities are more likely to be near these sources of pollution due to underlying structural inequalities, which again points to the plausibility of the shared confounding mechanism.
@@ 6.1 Factor loadings and implications for unmeasured confounding
Estimates of the factor model parameters are shown in Figure 3. The maximum number of factors that can be identified in factor analysis with 6 exposures and 6 outcomes (including the negative control outcome) is $m=3$ for both treatment and outcome models. We proceed with $m=3$ for this reason, as it will provide the most conservative inferences about causal effects of interest. We see that there are large factor loadings, particularly in the $B$ matrix due to high correlation among the exposures. We find that ozone has very small factor loadings, indicating little dependence with the other exposures. The first column of $\Gamma$ largely reflects correlation among pulmonary outcomes: COPD, lung cancer and prior COPD. As such, prior COPD will be an especially useful null control outcome for COPD and lung cancer. The factor loadings also provide us with information about $R^{2}_{a^{\top}Y\sim U|T,X}$ and $R^{2}_{d^{\top}T\sim U|X}$ for each outcome, $a$, and exposure, $d$ (see Table 1). Larger values of these R-squared values will generally lead to wider partial identification regions for those exposures or outcomes due to the increased potential for unmeasured confounding bias.
@@ 6.2 Causal effect of organic carbon and nitrates
By way of illustration, we explore the partial identification regions for the effect of both organic carbon and nitrates on the outcomes of interest. The results for other exposures are included in Appendix E. The partial identification regions for the other exposures generally contain zero and are wide for elemental carbon, sulfates, and ammonium due to the large partial R-squared values seen in Table 1. We incorporate two different constraints when calculating the partial identification regions for each exposure of interest: 1) using prior COPD as a negative control outcome, and 2) using prior COPD as a negative control outcome and placing an effect size constraint on the effect of that exposure on asthma hospitalizations with $z_{l}=-0.05$ and $z_{u}=0.2$. The lower bound stems from prior research and biologic implausibility of pollution having a protective effect on health, while the upper bound is based on prior literature looking at larger shifts in particulate matter than those considered here, which have estimated smaller effects than this upper bound (Di et al., 2017). Complex public health outcomes such as hospitalization rates are driven by a host of factors (smoking and obesity, among others), and particulate matter only explains a small part of the variability in these outcomes. Given that our outcomes are standardized, it is therefore implausible to have an effect size of 0.25 or 0.5 as this would imply that pollution explains a large portion of the variability in hospitalization rates. To ensure that any results we see are not strictly due to the asymmetric nature of our bound, we apply symmetric bounds in Appendix E and find similar conclusions.
@@ 7 Discussion
In this paper, we develop a framework for identifying partial identification regions in the presence of unmeasured confounding when the data contain multiple treatments and multiple outcomes. We illustrate how a factor confounding assumption can be leveraged to learn more about the range of plausible causal effects across multiple estimands. For one, we obtain bounds on causal effects that are identifiable from the observed data alone. These partial identification regions can be reduced in size significantly under reasonable constraints that are known a priori given subject-matter expertise. One of the constraints we consider are negative control assumptions, for which we develop theory and algorithms to produce refined partial identification regions. Alternatively, prior knowledge on plausible effect sizes or constraints on partial $R^{2}$ values obtained through calibration to observed covariates, can be used to provide more informative partial identification regions. Overall, our theoretical and empirical studies suggest that the proposed framework yields new insights about causal effects even in the presence of unmeasured confounding variables. This was evident in a nationwide analysis of the effects of particulate matter components on cause specific hospitalizations using Medicare claims data, which showed a strong, harmful effect of organic carbon on public health outcomes.
@@ Acknowledgements
Research described in this article was conducted under contract to the Health Effects Institute (HEI), an organization jointly funded by the United States Environmental Protection Agency (EPA) (Assistance Award No. CR-83590201) and certain motor vehicle and engine manufacturers. The contents of this article do not necessarily reflect the views of HEI, or its sponsors, nor do they necessarily reflect the views and policies of the EPA or motor vehicle and engine manufacturers.
@@ Appendix A Technical Details
A.1 Proof of Theorem 1
Proof.
@@ A.1 Proof of Theorem 1
A.2 Proof of Theorem 2
Proof.
We first show the identification of $\|a^{\top}\Gamma\|_{2}$. Under the condition (C1) of Assumption 4, $\Gamma\Gamma^{\top}$ and $\Sigma_{y|t,u}$ is uniquely determined from
@@ A.2 Proof of Theorem 2
by applying Lemma 5.1 and Theorem 5.1 of Anderson & Rubin (1956) to the factor model (3), $\Gamma$ is identified up to rotations from the right under the factor confounding assumption. In other words, $\Gamma$ is identified up to multiplication on the right by an orthogonal matrix, and thus any admissible value for $\Gamma$ can be written as $\widetilde{\Gamma}=\Gamma R$ with an arbitrary $m\times m$ orthogonal matrix $R$. Then $\|a^{\top}\Gamma\|_{2}$, which is rotation-invariant, is identified as follows.
@@
Next, we show the identification of $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$, by applying the same idea to the factor model (2). Under condition (C2) of Assumption 4, $BB^{\top}$ and $\Sigma_{t|u}$ are uniquely determined from $\text{Cov}(T)=BB^{\top}+\text{Cov}(\epsilon_{t})=BB^{\top}+\Sigma_{t|u}$, and any admissible value for $B$ can be written as $\widetilde{B}=BR$ where $R$ is again an arbitrary $m\times m$ orthogonal matrix. Then, the conditional distribution of $U$ given $T$ with $\widetilde{B}$ is
@@
$=R^{\top}\Sigma_{u|t}R,$
where $\mu_{u|t}$ and $\Sigma_{u|t}$ are from (4). Then, we can show that
@@
where $RR^{\top}=I$, which implies that $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$ is identified. Combining the identification results, the bias bound is identifiable under the conditions (C1) and (C2).
@@
A.3 Proof of Theorem 3
Proof.
We assume that we only have a single negative control contrast. This provides us information about the plausible values of $b_{j}^{\top}\Gamma$. To ensure that the negative control assumptions are compatible, the solution for (13) exists as
@@ A.3 Proof of Theorem 3
holds (Penrose 1955), where $M_{j}^{\dagger}$ denotes a generalized inverse of $M_{j}$, $M_{j}M_{j}^{\dagger}$ is the projection matrix onto the column space of $M_{j}$, and $M_{j}^{\dagger}M_{j}$ is the projection matrix onto the row space of $M_{j}$. Note here that if $c_{j}=m$, we have that $M_{j}^{\dagger}$ is a standard inverse instead of a generalized inverse, and we are able to identify $b_{j}^{\top}\Gamma$.
@@
$=\underbrace{\frac{k_{ab}}{k_{bb}}b_{j}^{\top}\check{\mathcal{G}}_{j}M_{j}^{\dagger}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(I)\text{ bias correction}}+\underbrace{\frac{k_{ab}}{k_{bb}}w_{b,j}(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(II)}$
$~{}~{}~{}+\underbrace{w_{a}\{I-({D}^{\top})({D}^{\top})^{\dagger}\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(III)}$
@@
$|(II)|$ $\leq\left\|\frac{k_{ab}}{k_{bb}}w_{b,j}(I-M_{j}M_{j}^{\dagger})\right\|_{2}\left\|(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\right\|_{2}$
@@
A.4 Extension to multiple negative controls
Theorem 5.
$K^{*}D^{*}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\quad\pm$
Proof.
Suppose that there are $J$ outcomes with at least one negative control treatment. Letting $\mathcal{Q}\subset\{1,\cdots,q\}$ denote a set of the indices of such outcomes and $\text{card}(\cdot)$ denote the cardinality of a set, we have that $J=\text{card}(\mathcal{Q})$. The negative control contrast provides us information about the plausible values of $b_{j}^{\top}\Gamma$ for $j\in\mathcal{Q}$. In particular, we can see the following. By Theorem 2 of Penrose (1955), a necessary and sufficient condition for (13) to have a solution is
in which case the general solution exists as
Under our factor modeling assumptions, we are able to identify
$a^{\top}\Gamma\Gamma^{\top}\boldsymbol{b}:=K_{ab}$ (23)
for some arbitrary $m$-dimensional row vector $w_{a}$. This follows
The bias for the estimand of interest is given by
$~{}~{}~{}~{}a^{\top}\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$
$=\underbrace{K_{ab}K_{bb}^{-1}\boldsymbol{D}^{(1)}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(I)\text{ bias correction}}+\underbrace{K_{ab}(\boldsymbol{D}\boldsymbol{D}^{\top})^{-1}\boldsymbol{D}^{(2)}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(II)}+\underbrace{w_{a}\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{(III)},$
$(II)=\sum_{j=1}^{J}\alpha_{j}w_{b,j}(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t},$
$|(II)|$ $\leq\sum_{j=1}^{J}|\alpha_{j}|\|w_{b,j}(I-M_{j}M_{j}^{\dagger})\|_{2}\|(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$
$|(III)|$ $\leq\left\|w_{a}\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\right\|_{2}\left\|\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\right\|_{2}$
which is equivalent to
$K^{*}\boldsymbol{D}^{*}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\quad\pm$
where $K^{*}:=K_{ab}K_{bb}^{-1}$ and $\boldsymbol{D}^{*}:=\boldsymbol{D}^{(1)}$. ∎
A.5 Proof of Theorem 4
@@
which is identifiable from our factor confounding assumptions.

### v3

@@ 5.3.2 Identification conditions with negative controls
While the focus of our paper is generally on partial identification, we show here that under certain scenarios, negative controls can point identify the causal effect of interest. This further shows the potential benefit of negative controls and provides connections between our work and identification scenarios in the literature on double negative controls (Miao et al.,, 2018; Shi et al., 2020a, ; Shi et al., 2020b, ; Hu et al.,, 2023).
@@
A proof of this result can be found in Appendix A. While these conditions may seem abstract, they do hold in some plausible scenarios. For example, scenario 1 holds automatically when we have $m$ negative control exposures, i.e., negative control pairs with the same outcome as the estimand of interest, but different exposure contrasts. Identification can also be obtained if additional assumptions are made, such as the existence of a so-called confounding bridge function (Miao et al.,, 2018), though we point readers to the literature on double negative controls for a discussion of such assumptions. Lastly, it is important to emphasize that within our framework, negative controls are still useful even if these conditions do not hold, as they can reduce the widths of the partial identification region substantially compared with the widths without negative controls. This is related to recent work on partial identification in the proximal causal inference literature where partial identification can be obtained when assumptions required for point identification do not hold (Ghassami et al.,, 2023).
@@ 6 Simulation study
Here, we present a brief simulation study to highlight the benefits of characterizing the joint partial identification region for multiple estimands. For simplicity, we focus on a single data set with sample size $n=10^{6}$ to effectively eliminate sampling variability and focus on the uncertainty caused by the potential presence of unmeasured confounders. For brevity, we leave explicit details of the data generation process to Appendix D. To summarize, we consider a problem with $k=10$ exposures and $q=6$ outcomes, with $m=3$ unmeasured confounders. The true effect sizes for all estimands are between $-1$ and 1, and there are a number of treatments with no effect on specific outcomes, leading to multiple potential negative controls. We use linear models throughout, which contains the true data generating mechanism, though similar results hold for more complex estimators of $g(\cdot)$. Our inferential goal is to estimate the effect of a particular exposure on a single outcome. We scaled the confounding and effect sizes so that the estimate obtained assuming no unmeasured confounding is 0, while the true value of the estimand is 1. To illustrate the different assumptions or constraints that one can place, we derive the partial identification region for this estimand under a variety of different assumptions, summarized below.
@@
Now we return to the motivating example and data analysis, which was introduced in Section 2. We standardize all exposures and outcomes to have mean zero and standard deviation one prior to running the analysis so that the effect estimates are all on the same scale across outcomes and exposure combinations. For example, an estimate of 0.1 for a particular shift in exposures would be interpreted as being expected to change the outcome by 0.1 outcome standard deviations. The estimands we look at are the effects of shifting one exposure from their 1st to 3rd quartile, while holding the remaining exposures fixed at their median value. We estimate $g(\cdot)$ and therefore the effects of the exposures (and confounders) on the outcome using a linear regression model, and use the aforementioned approaches to assess the robustness of these estimates to unmeasured confounding bias. For more details on the estimation of all model parameters, including factor model parameters, see Appendix C.
@@
We also note that the factor confounding assumption (Assumption 4) is plausible in this application. All outcomes correspond to hospitalization rates for different diseases, and it is reasonable that certain characteristics, which are not measured in our observed covariates, are broadly associated with poor health outcomes. For instance, these unmeasured characteristics likely reflect features like health-seeking behavior or proximity to and/or quality of health care access. Similarly, exposures tend to be elevated in areas with various pollutant emitting sources such as roadways, power plants, and factories. Certain communities are more likely to be near these sources of pollution due to underlying structural inequalities, which again points to the plausibility of the shared confounding mechanism.
@@ 7.1 Factor loadings and implications for unmeasured confounding
Estimates of the factor model parameters are shown in Figure 3. The maximum number of factors that can be identified in factor analysis with 6 exposures and 6 outcomes (including the negative control outcome) is $m=3$ for both treatment and outcome models. We proceed with $m=3$ for this reason, as it will provide the most conservative inferences about causal effects of interest. We see that there are many large factor loadings, particularly in the exposure loading matrix, B. However, we find that ozone has very small factor loadings, indicating little dependence with the other exposures. The first column of the outcome loading matrix, $\Gamma$, largely reflects correlation among pulmonary outcomes: COPD, lung cancer and prior COPD. As such, prior COPD is an especially useful null control outcome for COPD and lung cancer. The factor loadings also provide information about $R^{2}_{a^{\top}Y\sim U|T,X}$ and $R^{2}_{d^{\top}T\sim U|X}$ for each outcome, $a$, and exposure, $d$ (see Table 1). Larger values of these R-squared values imply wider partial identification regions for those exposures or outcomes due to the increased potential for unmeasured confounding bias.
@@ 7.2 Causal effect of organic carbon and nitrates
We now explore the partial identification regions for the effect of both organic carbon and nitrates on the outcomes of interest. We focus on these two exposures in the manuscript for simplicity, and because these two exposures illustrate two distinct scenarios where either the partial identification regions are informative about the causal effect, or they are not informative. The results for other exposures are included in Appendix E. The partial identification regions for the other exposures generally contain zero and are wide for elemental carbon, sulfates, and ammonium due to the large partial R-squared values seen in Table 1. We incorporate two different constraints when calculating the partial identification regions for each exposure of interest: 1) using prior COPD as a negative control outcome, and 2) using prior COPD as a negative control outcome and placing an effect size constraint on the effect of that exposure on asthma hospitalizations with $z_{l}=-0.05$ and $z_{u}=0.2$. The lower bound stems from prior research and biologic implausibility of pollution having a protective effect on health, while the upper bound is based on prior literature looking at larger shifts in particulate matter than those considered here, which have estimated smaller effects than this upper bound (Di et al.,, 2017). Complex public health outcomes such as hospitalization rates are driven by a host of factors (smoking and obesity, among others), and particulate matter only explains a small part of the variability in these outcomes. Given that our outcomes are standardized, it is therefore implausible to have an effect size of 0.25 or 0.5 as this would imply that pollution explains a large portion of the variability in hospitalization rates. To ensure that any results we see are not strictly due to the asymmetric nature of our bound, we apply symmetric bounds in Appendix E and find similar conclusions.
@@ 8 Discussion
In this paper, we develop a framework for identifying partial identification regions in the presence of unmeasured confounding when the data contain multiple treatments and multiple outcomes, which is particularly useful for studying the health effects of multivariate air pollutants. We illustrate how a factor confounding assumption can be leveraged to learn more about the range of plausible causal effects across multiple estimands. For one, we obtain bounds on causal effects that are identifiable from the observed data alone. These partial identification regions can be reduced in size significantly under reasonable constraints that are known a priori given subject-matter expertise, which one typically has in environmental studies. One of the constraints we consider are negative control assumptions, for which we develop theory and algorithms to produce refined partial identification regions. Alternatively, prior knowledge on plausible effect sizes or constraints on partial $R^{2}$ values obtained through calibration to observed covariates, can be used to provide more informative partial identification regions. Overall, our theoretical and empirical studies suggest that the proposed framework yields new insights about causal effects even in the presence of unmeasured confounding variables. This was evident in a nationwide analysis of the effects of particulate matter components on cause specific hospitalizations using Medicare claims data, which showed a strong, harmful effect of organic carbon on public health outcomes.
@@ Acknowledgements
Research described in this article was conducted under contract to the Health Effects Institute (HEI), an organization jointly funded by the United States Environmental Protection Agency (EPA) (Assistance Award No. CR-83590201) and certain motor vehicle and engine manufacturers. The contents of this article do not necessarily reflect the views of HEI, or its sponsors, nor do they necessarily reflect the views and policies of the EPA or motor vehicle and engine manufacturers. The computations in this paper were run on the FASRC Cannon/FASSE cluster supported by the FAS Division of Science Research Computing Group at Harvard University.
@@ Appendix A Technical Details
A.1 Proofs of Proposition 1, and Theorems 1, 2, 3, and 4
Proof of Theorem 1.
@@ A.1 Proofs of Proposition 1, and Theorems 1, 2, 3, and 4
Proof of Theorem 2.
We first show the identification of $\|a^{\top}\Gamma\|_{2}$. Under the first condition of Assumption 4, $\Gamma\Gamma^{\top}$ and $\Sigma_{y|t,u}$ is uniquely determined from
@@
by applying Lemma 5.1 and Theorem 5.1 of Anderson and Rubin, (1956), $\Gamma$ is identified up to rotations from the right under the factor confounding assumption.
In other words, $\Gamma$ is identified up to multiplication on the right by an orthogonal matrix, and thus any admissible value for $\Gamma$ can be written as $\widetilde{\Gamma}=\Gamma R$ with an arbitrary $m\times m$ orthogonal matrix $R$. Then $\|a^{\top}\Gamma\|_{2}$, which is rotation-invariant, is identified as follows.
@@
Next, we show the identification of $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$, by applying the same idea to the treatment model. Under condition 2 of Assumption 4, $BB^{\top}$ and $\Sigma_{t|u}$ are uniquely determined from $\text{Cov}(T)=BB^{\top}+\text{Cov}(\epsilon_{t})=BB^{\top}+\Sigma_{t|u}$, and any admissible value for $B$ can be written as $\widetilde{B}=BR$ where $R$ is again an arbitrary $m\times m$ orthogonal matrix. Then, the conditional distribution of $U$ given $T$ with $\widetilde{B}$ is
@@
$=R^{\top}\Sigma_{u|t}R.$
Then, we can show that
@@
where $RR^{\top}=I$, which implies that $\|\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$ is identified. Combining the identification results, the bias bound is identifiable under the conditions of Assumption 4.
@@
Proof of Proposition 1 and Theorem 3.
Here we focus on Theorem 3, but the result in Proposition 1 can be found as a special case of the result in Theorem 3 and therefore we do not provide an explicit proof. We assume that we only have a single negative control contrast. This provides us information about the plausible values of $b_{j}^{\top}\Gamma$. To ensure that the negative control assumptions are compatible, the solution for (16) exists as
@@
holds (Penrose,, 1955), where $M_{j}^{\dagger}$ denotes a generalized inverse of $M_{j}$, $M_{j}M_{j}^{\dagger}$ is the projection matrix onto the column space of $M_{j}$, and $M_{j}^{\dagger}M_{j}$ is the projection matrix onto the row space of $M_{j}$. Note here that if $c_{j}=m$, we have that $M_{j}^{\dagger}$ is a standard inverse instead of a generalized inverse, and we are able to identify $b_{j}^{\top}\Gamma$.
@@
$=\underbrace{\frac{k_{ab}}{k_{bb}}b_{j}^{\top}\check{\mathcal{G}}_{j}M_{j}^{\dagger}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{ (I) bias correction}}+\underbrace{\frac{k_{ab}}{k_{bb}}w_{b,j}(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{(II)}}$
$~{}~{}~{}+\underbrace{w_{a}\{I-({D}^{\top})({D}^{\top})^{\dagger}\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{(III)}}$
@@
[…]
@@
Proof of Theorem 4.
@@
which is identifiable from our factor confounding assumptions. ∎
A.2 Extension to multiple negative controls
Theorem 5.
$K^{*}D^{*}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\quad\pm$
Proof of Theorem 5.
Suppose that there are $J$ outcomes with at least one negative control treatment. Letting $\mathcal{Q}\subset\{1,\cdots,q\}$ denote a set of the indices of such outcomes and $\text{card}(\cdot)$ denote the cardinality of a set, we have that $J=\text{card}(\mathcal{Q})$. The negative control contrast provides us information about the plausible values of $b_{j}^{\top}\Gamma$ for $j\in\mathcal{Q}$. In particular, we can see the following. By Theorem 2 of Penrose, (1955), a necessary and sufficient condition for (16) to have a solution is
in which case the general solution exists as
Under our factor modeling assumptions, we are able to identify
$a^{\top}\Gamma\Gamma^{\top}\boldsymbol{b}:=K_{ab}$ (28)
for some arbitrary $m$-dimensional row vector $w_{a}$. This follows
The bias for the estimand of interest is given by
$~{}~{}~{}~{}a^{\top}\Gamma\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}$
$=\underbrace{K_{ab}K_{bb}^{-1}\boldsymbol{D}^{(1)}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{ (I) bias correction}}+\underbrace{K_{ab}(\boldsymbol{D}\boldsymbol{D}^{\top})^{-1}\boldsymbol{D}^{(2)}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{(II)}}+\underbrace{w_{a}\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}}_{\text{(III)}},$
[…]
$|(\text{II})|$ $\leq\sum_{j=1}^{J}|\alpha_{j}|\|w_{b,j}(I-M_{j}M_{j}^{\dagger})\|_{2}\|(I-M_{j}M_{j}^{\dagger})\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\|_{2}$
$|(\text{III})|$ $\leq\left\|w_{a}\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\right\|_{2}\left\|\left\{I-(\boldsymbol{D}^{\top})(\boldsymbol{D}^{\top})^{\dagger}\right\}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\right\|_{2}$
which is equivalent to
$K^{*}\boldsymbol{D}^{*}\Sigma_{u|t}^{-1/2}\mu_{u|\Delta t}\quad\pm$
where $K^{*}:=K_{ab}K_{bb}^{-1}$ and $\boldsymbol{D}^{*}:=\boldsymbol{D}^{(1)}$. ∎